OpenStudy (anonymous):

How do I use the trapezoidal rule to approximate the area bound by y= sin^2x, x=0, y=0, and x=1? Help appreciated! ^-^

OpenStudy (anonymous):

oh and I'm supposed to use n= 4 so I am thinking the integral is from 0 to 1 right? so since I need 4 segments, should I use .25, .5, .75, and 1? I tried the problem once this way and got .316 but I am pretty sure this is incorrect. Anyways, I'm stuck, so I would love some help!

OpenStudy (anonymous):

The trapezoidal rule is that the area at every segment from a to b can be approximated by \(Y(a, b) = (b-a)\frac{f(a) + f(b)}{2}\) (nonstandard symbol \(Y\)). In your case, the approximation becomes \(Y = Y(0, .25) + Y(.25, .5) + Y(.5 + .75) + Y(.75, 1)\). Because \((b-a)\) is constant throughout all four terms, the expression can be simplified using the identity \(a(b+c)=ab+ac\): \(Y = \frac{b - a}{2}\left(\frac{f(0)}{2}) + f(.25) + f(.5) + f(.75) + \frac{f(1)}{2}\right)\) where \(b - a = .25\).

OpenStudy (anonymous):

i got 0.27743

OpenStudy (anonymous):

lets see what i get

OpenStudy (anonymous):

thank you for your replies! I'll just take a minute to look them over and then see if I made any mistakes.

OpenStudy (anonymous):